Parramore Corp has $11 million of sales, $1 million of inventories, $4 million of receivables, and $2 million of payables. Its cost of goods sold is 75% of sales, and it finances working capital with bank loans at an 8% rate. Assume 365 days in year for your calculations. What is Parramore's cash conversion cycle (CCC)? Do not round intermediate calculations. Round your answer to two decimal places. days If Parramore could lower its inventories and receivables by 7% each and increase its payables by 7%, all without affecting sales or cost of goods sold, what would be the new CCC? Do not round intermediate calculations. Round your answer to two decimal places. days How much cash would be freed up, if Parramore could lower its inventories and receivables by 7% each and increase its payables by 7%, all without affecting sales or cost of goods sold? Write out your answer completely. For Example, 13.2 million should be entered as 13,200,000. Do not round intermediate calculations. Round your answer to the nearest dollar. $ By how much would pretax profits change, if Parramore could lower its inventories and receivables by 7% each and increase its payables by 7%, all without affecting sales or cost of goods sold? Write out your answer completely. For Example, 13.2 million should be entered as 13,200,000. Do not round intermediate calculations. Round your answer to the nearest dollar. $
Answer #1
1] Existing New
Sales $ 1,10,00,000 $   1,10,00,000
COGs $         82,50,000 $      82,50,000
Inventories $         10,00,000 $   9,30,000.00
Receivables $         40,00,000 $ 37,20,000.00
Payables $         20,00,000 $ 21,40,000.00
Days' inventory outstanding [DIO] [Inventory*365/COGS] 44.24 41.15
Days' sales outstanding [DSO] [Receivables*365/Sales] 132.73 123.44
Days payables outstanding [DSO] [Payables*365/COGS] 88.48 94.68
Cash conversion cycle [CCC] = DIO+DSO*DPO = 88.48 69.90
2] NWC $         30,00,000 $ 25,10,000
Cash freed up = 3000000-2510000 = $         4,90,000
3] Change in pre tax profits = 490000*8% = $            39,200